An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to
get a 12% mixture?

Answer by CharlesG2(834)

An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to
get a 12% mixture?

%A * amount A + %B * amount B = %C * amount C
15% * 20 + 10% * B = 12% * (20 + B)
0.15 * 20 + 0.1 * B = 0.12 * (20 + B)
3 + 0.1B = 2.4 + 0.12B
0.6 = 0.12B - 0.1B
0.6 = 0.02B
B = 30
30 pounds of the second alloy to get 50 pounds of 12% tin